Suppose that a batch of parts is played with copper in an electrolytic bath running at 0.15V and 15A for exactly 2 hrs. What is the energy cost of this process if the electric utility charges the company Php. 4.05/KW-hr.
1
Expert's answer
2020-11-14T13:52:33-0500

Energy = VIt

Time t in seconds = 2hr which is equal to 7200sec

Energy = 0.15×15×7200


Energy = 16200 joules = 16.2 KW-hr

Energy cost = 16.2×4.05 =65.61